A diffraction grating produces a second-order principal maximum at 50.6° to the normal when being illuminated normally with light of wavelength 644nm. Calculate the number of lines per millimeter of the grating and the total number of principal maxima produced.


1
Expert's answer
2021-02-18T18:46:04-0500

Given Parameters are:-

order, n=2

Angle, "\theta=50.6^{\circ}"

Wavelength, "\lambda=644nm=644\times 10^{-6}\text{mm}"


As we know,

"n\lambda=dsin\theta"

"2\times 644\times 10^{-9}=d\times sin50.6^{\circ}"


"\Rightarrow d=\dfrac{1288\times 10^{-6}}{sin50.6^{\circ}}"


="1666.8\times 10^{-6}=1.67\times10^{-3}mm"


So, The number of lines per mm is "\dfrac{1}{d}=\dfrac{1}{1.67\times 10^{-3}}=598.8"


Hence The number of lines per mm is "598.8\equiv599"


As there are second order principal maxima so,


Total of "5" images is seen on screen from which 3 are principal maxima.


Total number of maxima produced"= 3"